OpenStudy (stevetron):

Order the group of quadratic functions from widest to narrowest graph. y=-9x^2, y=1\3x^2, y=-1\2x^2

OpenStudy (anonymous):

Y=-9x^2 Y=1/2x^2 Y=1/3x^2

OpenStudy (stevetron):

thank you